Tavon has a gift card for $80 that loses $3.50 for each​ 30-day period it is not used. He has another gift card for $70 that loses $3 for each​ 30-day period it is not used.
1
Expert's answer
2020-09-17T14:03:42-0400

80 - 3.50x ........ (i)

70 - 3x ..............(ii)

Set them equal;

80 - 3.50x = 70 - 3x

Solve for x. 80-70 =3.50x - 3x

10=0.50x

20=x

There will be 20 days out of the 30-day period until the cards will be equal in value.